Find the values of n for which the equation (n-1)²uxx - y^2n uyy = ny^(2n-1)uy is
i) parabolic ii) hyperbolic.
1
Expert's answer
2020-03-19T15:56:10-0400

"(n-1)²u_{xx} - y^{2n} u_{yy} = ny^{2n-1}u_y\\
A(x,y)u_{xx}+2B(x,y)u_{xy}+C(x,y)u_{yy}=\\
=F(x,y,u,u_x,u_y)\\"

The type of second-order PDE at a point 

"(x_0,y_0)"  depends on the sign of the discriminant defined as

"\Delta(x_0,y_0)=B^2-AC"

In our case

"A=(n-1)^2, B=0, C=-y^{2n}\\
\Delta =0^2 -(n-1)^2(-y^{2n})=(n-1)^2y^{2n}"

i) parabolic

"\Delta(x_0,y_0)=0\implies\\
(n-1)^2y^{2n}=0\implies\\
n=1 \quad or \quad y=0."

ii) hyperbolic

"\Delta(x_0,y_0)>0\implies\\
(n-1)^2y^{2n}>0\implies\\
n\neq1 \quad \quad y\neq 0."